How much heat required to convert 36 grams (2 moles) of liquid water at 4.0 oC to liquid water at 50 oC? The delta-H of fusion for H2O is 6.02 kJ/mol and the heat capacity of liquid water is 4.18 J/goC

A) 36 kJ

B) 12.04 kJ

C) 8.36 kJ

D) 4.18 kJ

E) 6.9 kJ
Answer #1

Here, Neight of liquid water - 36 gm. 2,млн +4°C -5 че от ѕве ОТ = Here Ту-т, To = soc т = 4с T-T, so t = 4с 4-6 с AT = AT =